OpenStudy (anonymous):

(w + x)(w + x) as an answer i got xw+x^2?

OpenStudy (id21):

it is the same as \[(w+x)^2=w^2+2wx+x^2\]

OpenStudy (anonymous):

w^2+2wx+x^2 foil method first outside inside last

OpenStudy (anonymous):

\[(w+x)(w+x)=w\times w + w\times x +x \times w + x \times x\]

OpenStudy (anonymous):

giving \[w^2+2wx+x^2\]

OpenStudy (anonymous):

w^2+wx+xw+x^2 w^2+2wx+x^2

OpenStudy (anonymous):

\[w ^2+2wx+x ^2\]